OpenStudy (anonymous):

find the critical numbers, g(t)=t*sqrt(1-t), t<1 pleas ehelp me on this one guys, and i forgot the last perosn who helped me but thank yous so much..my comp froze:S

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

g'(t) = sqrt(1-t) - t/2sqrt(1-t) We'll put g'(t) = 0 sqrt(1-t) - t/2sqrt(1-t) = 0 2(1-t) - t = 0 2 - 2t - t = 0 -3t = -2 t = 2/3 critical number t = 2/3

OpenStudy (anonymous):

thank you so much!!!:)

OpenStudy (anonymous):

\[g'(t)=(2-3t)/2(1-t)^{1/2}\] critical values are \[2-3t =0\] and \[2(t-1)^{1/2}=0\] thus, t= 1 and t=2/3